How many liters of hydrogen at STP can be produced from 28.0 g of zinc reacting with H2SO4?
1
Expert's answer
2020-04-24T13:34:00-0400

Solution:

1 mol of Zn has mass of 65.38 g.

The amount of Zn is (28.0 g) / (65.38 g/mol) = 0.428 mol.

The balanced chemical equation:

Zn + H2SO4 = ZnSO4 + H2

According to the chemical equation: n(Zn) = n(H2).

The amount of H2 produced is the same as the amount of Zn consumed (0.428 mol).

n(H2) = 0.428 mol.


1 mol of ideal gas will occupy 22.4 L at STP.

The H2 will occupy 0.428 mol × 22.4 L/mol = 9.5872 L = 9.59 L.

V(H2) = 9.59 L.


Answer: 9.59 L of hydrogen (H2) can be produced.